OpenStudy (anonymous):

Find the exact length of the curve. x = (y^4/8)+(1/4y^2) 1 ≤ y ≤ 2

OpenStudy (phi):

if the problem is \[ x = \frac{y^4}{8} + \frac{1}{4y^2} \] first find dx/dy

OpenStudy (phi):

then evaluate \[ \int_1^2 \sqrt{1+ \left( \frac{dx}{dy}\right)^2} \ dy \]